The sum of two numbers is twenty five and the sum of their square is three hundred and twenty five. Find the numbers.
Answer by stanbon(48568)

The sum of two numbers is twenty five and the sum of their square is three hundred and twenty five. Find the numbers.
-------------
Let one of the numbers be "x"; the other must be "25-x".
EQUATION:
x^2+(25-x)^2=325
x^2+625-50x+x^2=325
2x^2-50x+300=0
x^2-25x+150=0
x=[25+-sqrt(625-4*150)]/2
x=[25+-5]/2
x=15or x=10
One of the numbers is 15; the other is 10
===========
Cheers,
Stan H.